Session: Mar 2009 (session 1)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Shanghai (Bo Xia at Sun Mar 1 06:33:19 2009) Minor failures :
We lost the last scan.

Onsala (Michael Lindqvist at Sun Mar 1 08:14:53 2009) Success :
One bad scan_check. 2009.060.00:28:00.30/scan_check/9,ew012b_on_no0055,mark4,64,2009y60d0h21m57.1163s,?,?,? Otherwise no known problems.

Medicina (Morsiani at Sun Mar 1 08:54:33 2009) Minor failures : RECO
lost scan from 06:20 (scan No0126) to the end of the schedule. Problems with one Hard Disc in MK5 module OSOD-070.

Westerbork (Geert Kuper at Sun Mar 1 15:22:39 2009) Success :
Wb used in phased array mode. No known problems

Torun (Kaz Borkowski at Mon Mar 2 10:30:22 2009) Minor failures : RECO
Nearly 3 hours (till 22:44 UT) in the beginning could not be recorded due to failure of Mk5A to write on nonempty DP in bank/bay A. Somewhat similar, but more severe (Magda had to come from Torun to help), situation as described in the feedback to GC031C. This problem is not related to a failing disk for the suspect DP can be used in bank B.

Effelsberg (Uwe Bach at Mon Mar 2 10:47:38 2009) Success :

Urumqi (Wenjun Yang at Mon Mar 2 10:55:55 2009) Success :

Noto (P. Cassaro at Fri Mar 6 15:44:29 2009) Minor failures : RECO
From 23:21 to 23:50 UT recording problems.

Jodrell_Bank (Alastair Gunn at Tue Mar 24 11:08:51 2009) Success :
The MkII telescope was used (Jb2). No known problems.

Cambridge (Alastair Gunn at Tue Mar 24 11:09:16 2009) Success :
Knockin telescope also recorded. No known problems.

Hartebeesthoek (dasrat at Sun Jul 5 04:30:21 2009) Success :